OpenStudy (anonymous):

Given the exponential function f(x)=2^x, find the range when the transformation is applied: g(x)=5(2+6f(23-x)). What would I do with the 5 in g(x)? Distributive property???

OpenStudy (anonymous):

You're correct on that

OpenStudy (anonymous):

How would I do it though?

OpenStudy (anonymous):

would it be g(x) = 10 + 6f(23-x) ?

OpenStudy (anonymous):

There are two functions. and F(x) is a function of G(x). What do you think the domain of 2^x is?

OpenStudy (anonymous):

XER

OpenStudy (anonymous):

Good. Now what is f(23-x)?

OpenStudy (anonymous):

same

OpenStudy (anonymous):

I meant what is it, not its domain? what does it mean?

OpenStudy (anonymous):

If f(x) = 2^x f(23-x) = ?

OpenStudy (anonymous):

f(-(x-23)) reflection in y-axis horizontal shift of 23 units to right

OpenStudy (anonymous):

I was hoping you would say its 2^(23-x)

OpenStudy (anonymous):

oh, i thought u meant what it meant as in transformations. But i get that. I just don't understand what to do with that 5 in front of g(x)

OpenStudy (anonymous):

now we can write the function G(x) as G(x) = 5(2+6(2^(23-x))

OpenStudy (anonymous):

can you see the substitution I made?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now distribute the 5.

OpenStudy (anonymous):

g(x) = 10+30(2^(23-x)

OpenStudy (anonymous):

what about 2^(23-x)? Does that get affected?

OpenStudy (anonymous):

Does this make more sense now? you can even distribute the 2^(23-x). I believe it becomes 2^23-2^x

OpenStudy (anonymous):

Oh ok so the transformed function becomes g(x) = 30(2^23-x) + 10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, thank you veryyy much :D

OpenStudy (anonymous):

your welcome =)